# algorithmic

### v2

@@ Abstract
A multi-event survival model predicts the time until an instance experiences each of several different events, given the instance’s description. Unlike competing-event models, the events here are not mutually exclusive and often exhibit statistical dependencies. Existing approaches for multi-event survival analysis have generally been limited, most focusing on producing simple risk scores for each event, rather than the time-to-event itself. To overcome these issues, we introduce MENSA, a novel deep learning approach for multi-event survival analysis. MENSA jointly learns representations of the input features while capturing the complex dependence structure among events. In practice, it attempts to optimize the sum of the traditional negative log-likelihood across events and a novel trajectory-based likelihood, which encourages the model to learn the temporal order in which events occur. Experiments on real-world clinical datasets show that MENSA consistently gives good discrimination performances and accurate time-to-event predictions in single-event, competing-risk, and multi-event problems. Additionally, MENSA is more computationally-efficient, requiring fewer parameters and FLOPs than multiple state-of-the-art survival baselines when applied to large-dimensional datasets (more than 100 features).
Keywords: Machine learning, survival analysis, competing risks, multi-event, precision medicine
@@ 1 Introduction
Many survival analysis models estimate the probability that some specific event (typically death) will occur for an individual at time $T$ later than $t$, for all times $t>0$ (Gareth et al., 2021, Ch. 11). Learning such models is challenging for standard regression tasks, as the training dataset can include many censored instances – cases where we only have a lower bound on the actual time of the occurrence. This problem can be extended to deal with many possible events – such as death from cancer, and also death from heart failure. This can involve competing risks, where the occurrence of one event excludes the others (e.g., one can either die from cancer or from heart failure, but not both) or multi-event scenarios, where the occurrence of one event does not exclude the others, such as different types of medical emergencies in hospitalized patients, e.g., acute respiratory failure (ARF) or shock (Tjandra et al., 2021), or adverse effects in patients who underwent a bone marrow transplantation, e.g., relapse or death (Marta Fiocco, 2008).
As a practical example, consider a drug trial over $t\in[0,100]$ days, where some patient $i$ (e.g., Mr. Smith) with some description $\bm{x}^{(i)}$ can experience 4 side effects: nausea (A), fatigue (B), fever (C) or headache (D). In survival analysis, we want to predict the survival probability (i.e., the probability that the event occurs at $T$ later than some time $t$) for each of these possible side effects over a finite time horizon, e.g., from 0 to 100 days. We can address this problem as a multi-event survival problem and build a single model to estimate the so-called individual survival distribution (ISD) of each of these four events for Mr. Smith, which we denote as $\hat{S}^{(i)}$ (see Figure 1). Thus, these ISDs give the probability of each of these 4 events occurring after $t$ days post-baseline, for all $t>0$. We can also use these ISDs to estimate the probable time-to-event for Mr. Smith, for example, when the survival (event) curve intersects the dashed horizontal line at 50%, for each of the events.
@@
In the current literature, some practitioners - e.g.Armstrong et al. (2014); Solomon et al. (2017) – simply use many single-event models for these multi-event problems. Perhaps one Cox Proportional Hazards (CoxPH) model (Cox, 1972) for cancer, and another for heart failure by modeling each event separately. This approach has multiple issues: (1) Separate models fail to leverage the shared patterns or features (covariates) between different events, however, real-world events often follow temporal or causal patterns that influence or correlate with each other, e.g., illness usually occurs before death (Fix and Neyman, 1951). (2) Training a new model for each event requires significant computational resources, particularly for large datasets or when there are many events of interest. (3) As the number of events increases, it becomes cumbersome to manage multiple single-event models. In contrast, a single multi-event model scales more efficiently by handling all events within a single architecture. There are, however, very few effective multi-event survival models in the literature, and most only deal with risk scores – i.e., simply attempting to identify the order of the events, both for each single individual, and also between individuals. However, we often need models that provide not only good discriminative performances (Tjandra et al., 2021), but also accurate time-to-event estimates (Qi et al., 2023a).
@@
In this paper, we propose a Multi-Event Network for Survival Analysis (MENSA). MENSA is a novel and scalable deep learning model for multi-event survival analysis, that learns the $K$ event distributions jointly, where each event is a convex combination of Weibull distributions. This approach relaxes the proportional hazards assumption about the underlying marginal survival distribution for each event, which is a significant limitation for Cox-based methods (Cox, 1972; Katzman et al., 2018; Lillelund et al., 2024b). As in previous works (Lee et al., 2018; Tjandra et al., 2021), we adopt a shared feature (covariate) layer in our architecture that captures the latent feature representation common to all events. Moreover, beyond simply optimizing the negative log-likelihood (NLL) over all events (Tjandra et al., 2021), we propose optimizing the sum of the NLL and a novel trajectory-based likelihood, which encourages the model to match the known temporal orderings between events based on prior knowledge. For example, consider the illness-death scenario (Fix and Neyman, 1951) with two events “illness (A)” and “death (B)”: If both events are observed for the patient $i$, then clearly event A must occur before event B, and so we want to maximize the probability of surviving event B, at the time of event A – so if A happened at 1 year, we want the probability of surviving B at 1 year to be very high.
MENSA is inspired by the multi-state framework, originally proposed by Andersen and Keiding (2002), where an event can be considered a transition from one state to another – e.g., alive to dead, or healthy to sick. This framework enables us to model the event-free distribution (i.e., the transient state) explicitly as if it were another event and perform likelihood estimation in conjunction with the events of interest. This allows practitioners to predict whether a patient will remain event-free for the duration of the study, along with other events. In addition, the MENSA framework easily extends to the single-event and competing-risk scenarios, making it suitable for a wide range of applications.
@@
We provide experimental evidence to show that MENSA can accurately discriminate (i.e., rank) instances by risk and make time-to-event predictions that improve over several shallow and deep learning baselines. We show that MENSA strikes a good balance between various evaluation tasks, e.g., discrimination, prediction, and calibration, without sacrificing either. Furthermore, MENSA uses fewer parameters and floating point operations per second (FLOPs) in high-dimensional datasets than many existing neural network approaches. Ablation experiments demonstrate that applying a shared feature layer can improve predictive performances if conditional dependencies exist between events, and that the proposed trajectory likelihood improves the accuracy of the time-to-event estimates.
Our code is publicly available at https://github.com/thecml/mensa
@@ 2 Related work
Single-event analysis. Time-to-event outcomes have been extensively studied within the field of survival analysis, primarily by the statistical, medical, and reliability engineering communities (Chen, 2024). Traditional single-event models aim to learn the time until a single event occurs for an instance, for example, the time to death in amyotrophic lateral sclerosis (ALS) (Kuan et al., 2023), the time to fall in community-dwelling elderly (Lillelund et al., 2024a), or various other events in economic and social sciences. In practice, “survival prediction” does not necessarily refer to actual mortality. The challenge in survival prediction is that the event times can be partially observed (i.e., censored), providing only a lower bound on the true event time. Multiple censoring mechanisms may exist in practice, but in this work we will focus on the most common form: right-censoring, where we know the event occurs after a certain time, but not exactly when.
Figure 2: Illustration of right-censorship in survival datasets.
@@
In a medical study as shown in Figure 2, some patients may experience the event (1st and 2nd patient), but some may be lost to follow-up (3rd patient), withdraw from the study (4th patient), or remain alive at the end of the study (5th patient). For these patients, the exact event time is known only to exceed their last contact or the end of the study. A wide range of models have been proposed for this setting, including the nonparametric Kaplan-Meier (KM) (Kaplan and Meier, 1958) and Nelson-Aalen (Nelson, 1969; Aalen, 1978) population estimators, semiparametric proportional hazards models for individuals, e.g., CoxPH (Cox, 1972; Faraggi and Simon, 1995; Katzman et al., 2018), and parametric survival models, e.g., the Weibull time-to-event model (Prentice and Kalbfleisch, 1979), Random Survival Forest (Ishwaran et al., 2008) and Multi-Task Logistic Regression (MTLR) (Yu et al., 2011). Chen (2024) provide an excellent introduction to the field and describe the transition from classical time-to-event methods to modern deep learning approaches.
@@
Competing-risks analysis. In competing-risks survival analysis, there are multiple events of interest, e.g., for a patient, we consider the patient dying from lung cancer, or from pneumonia, as two different events, which are mutually-exclusive. We thus want to learn a model that can predict, for each individual, which event will occur first and the time until this earliest event. Concerning censoring in such scenarios, competing risks is a form of dependent censoring (Emura and Chen, 2018, Ch. 1), as one risk (event) dependently censors the other, and these risks are assumed to have some prior relationship. However, estimating competing-risks data is more challenging than single-event analysis due to the identifiability issue (Tsiatis, 1975); since events are mutually-exclusive, the first event will censor the other events, preventing direct observation of their dependence structure. Given these challenges, multiple methods have been proposed for competing-risks scenarios, usually assuming that events are conditionally independent given the features.
@@
Lee et al. (2018) proposed DeepHit, a discrete-time neural network of the form $f_{\theta}:\bm{x}\rightarrow[0,1]^{T_{\text{max}}\times K}$ with parameter $\theta$, number of events $K$, and some maximum observed time $T_{\text{max}}$. DeepHit contains a shared covariate layer and $K$ cause-specific sub-networks; its learning process learns $\theta$ by maximizing the log-likelihood of the joint distribution of $K$ competing events and a cause-specific ranking loss.
@@
Kim et al. (2021) proposed MTLR-CR, an extension of the discrete MTLR framework for competing risks. MTLR (Yu et al., 2011) is a neural network implementation of a discrete-time survival model that predicts the survival distribution using a sequence of dependent logistic regressors. This is then extended to competing-risks scenarios by learning sequences of logistic regressors for the number of discrete time times, for every $K$ event.
@@
Nagpal et al. (2021) proposed Deep Survival Machines (DSM), a parametric survival model based on a mixture of predefined primitive distributions, e.g., a mixture of Weibulls. DSM first learns a common feature representation of all competing risks, and then a $k$th-event distribution for each event using cause-specific maximum likelihood estimation (MLE), treating all events other than $k$ as censored. However, this approach does not explicitly model the density or survival function relationships between events on a one-to-one basis. As a result, it may not account for potential interdependencies between events.
@@
Multi-event analysis. In multi-event analysis, we again consider multiple critical events, but these are not necessarily mutually-exclusive, as a patient may experience none, one or any subset of events at different times. For example, Mr. Smith may experience the onset of lung cancer at 1 year and may also experience the onset of pneumonia at 1.3 years, meaning that he would have both events (onset of lung cancer and of pneumonia). Multi-event analysis goes back as early as 1951, when Fix and Neyman (1951) introduced the illness-death model, which has been used in multiple subsequent works (Sverdrup, 1965). Unlike competing-risks, multi-event analysis does not necessarily suffer from the identifiability issue (Tsiatis, 1975), as the events are not necessarily mutually-exclusive. However, predicting the outcome of several (possibly related) events instead of just one poses another challenge: When events are, in fact, competing, naive survival models that do not account for the dependency structure between events may produce biased estimates.
@@
Larson (1984) adopted the popular CoxPH model (Cox, 1972) for multi-event survival analysis by modeling each event separately – an idea that has gained popularity since then (Armstrong et al., 2014; Solomon et al., 2017), but which fails to capture any dependencies between the events. Andersen and Keiding (2002) proposed a multi-state model that learns a model for each event transition (e.g., given a person is diagnosed with lung cancer, it then predicts the probability of pneumonia), but this also fails to leverage shared covariate information among the events. Several works have addressed this restriction, e.g., using a shared frailty term across models (Jiang and Haneuse, 2017), applying forms of regularization (Li et al., 2016; Wang et al., 2017) or by modeling the joint survival distribution (Hsieh and Wang, 2018). However, these approaches require assumptions on the distribution of the frailty term, the proportional hazards assumption, or assumptions on the form of the joint survival distribution, respectively.
@@
Tjandra et al. (2021) proposed a hierarchical multi-event model that learns the probability of event occurrence at different hierarchical time scales, using coarse predictions to iteratively guide predictions at finer levels – for example, from monthly to daily predictions. Given a complex task over a long time span, the model decomposes this task into a sequence of simpler tasks that serve as intermediate guides. Thus, the hierarchical framework progressively predicts survival across multiple time granularities, starting with broader intervals (e.g., months) and refining down to finer ones (e.g., weeks, then days). Formally, their model is a multi-task network that takes $x$ as input and outputs $\hat{P}$, a set of probability distributions for each $k$ event. Event-specific subnetworks, $\phi_{k}$, estimate the probability distribution $\hat{p}_{k}$ hierarchically. At granularity $m$, the probability distribution $\hat{p}_{km}$ is constructed from $\psi_{km}(\Phi(x))$ and the previous granularity level $\hat{p}_{k,m-1}$, ensuring that predictions at finer time scales are guided by coarser ones. However, there are two primary limitations to this approach: (1) the hierarchical component scales poorly with larger time durations and has a significant parametric overhead. (2) the method is optimized solely for discriminative loss –- i.e., ranking using Harrell’s concordance index (CI) (Harrell Jr. et al., 1982). However, strong ranking performance does not necessarily translate into accurate time-to-event predictions (Qi et al., 2023a), and vice versa. In our experiments, we will highlight these two limitations to the hierarchical model, i.e., its computational overhead and lack of predictive accuracy. Lastly, in Tjandra et al. (2021), the model is not evaluated on any real-world, multi-event dataset, but only on synthetic data. This can be problematic because synthetic datasets may not capture the complexity, variability, and noise present in real-world data.

### v3

@@ Abstract
We introduce MENSA, a novel deep learning model for multi-event survival analysis, which predicts the time until an instance experiences multiple distinct events based on its features. MENSA learns a shared representation of the input features while capturing the complex dependence structures between events. In practice, it optimizes the sum of the traditional negative log-likelihood across events and a novel trajectory-based likelihood, which encourages the model to learn the temporal order in which events occur. Experiments on real-world clinical datasets demonstrate that MENSA improves risk and time-to-event prediction compared to state-of-the-art models across single-event, competing-risk, and multi-event settings. Moreover, MENSA achieves this with fewer parameters and lower computational cost (FLOPs) than several deep learning baselines, particularly in high-dimensional feature spaces (more than 100 features).
keywords:
Machine learning, survival analysis, competing risks, multi-event, precision medicine
@@ 1 Introduction
Many survival analysis models estimate the probability that some specific event (typically death) will occur for an individual at time $T$ later than $t$, for all times $t>0$ [13, Ch. 11]. Learning such models is challenging for standard regression tasks, as the training dataset can include many censored instances – cases where we only have a lower bound on the actual time of the occurrence. This problem can be extended to deal with many possible events – such as death from cancer, and also death from heart failure. This can involve competing risks [10], where the occurrence of one event excludes the others (e.g., one can either die from cancer or from heart failure, but not both) or multi-event [2], where the occurrence of one event does not exclude the others, such as different types of medical emergencies in hospitalized patients, e.g., acute respiratory failure (ARF) or shock [51], or adverse effects in patients who underwent a bone marrow transplantation, e.g., relapse or death [35].
As a practical example, consider a drug trial over $t\in[0,100]$ days, where some patient $i$ (e.g., Mr. Smith) with some description $\bm{x}^{(i)}$ can experience 4 side effects: nausea (A), fatigue (B), fever (C) or headache (D). In survival analysis, we want to predict the survival probability (i.e., the probability that the event occurs at $T$ later than some time $t$) for each of these possible side effects over a finite time horizon, e.g., from 0 to 100 days. We can address this problem as a multi-event survival problem and build a single model to estimate the so-called individual survival distribution (ISD) of each of these four events for Mr. Smith, which we denote as $\hat{\bm{S}}^{(i)}$ (see Figure 1). Thus, these ISDs give the probability of each of these 4 events occurring after $t$ days post-baseline, for all $t>0$. We can also use these ISDs to estimate the probable time-to-event for Mr. Smith, for example, when the survival (event) curve intersects the dashed horizontal line at 50%, for each of the events.
@@
In the current literature, some practitioners – e.g.Armstrong et al. [3], Solomon et al. [49] – simply use many single-event models for these multi-event problems. Perhaps one Cox Proportional Hazards (CoxPH) model [6] for cancer, and another for heart failure by modeling each event separately. This approach has multiple issues: (1) Separate models fail to leverage the shared patterns or features (covariates) between different events, however, real-world events often follow temporal or causal patterns that influence or correlate with each other, e.g., illness usually occurs before death [11]. (2) Training a new model for each event requires significant computational resources, particularly for large datasets or when there are many events of interest. (3) As the number of events increases, it becomes cumbersome to manage multiple single-event models. In contrast, a single multi-event model scales more efficiently by handling all events within a single architecture. There are, however, very few effective multi-event survival models in the literature, and most only deal with risk scores – i.e., simply attempting to identify the order of the events, both for each single individual, and also between individuals. However, we often need models that provide not only good discriminative performances [51], but also accurate time-to-event estimates [40].
@@
In this paper, we propose a Multi-Event Network for Survival Analysis (MENSA). MENSA is a novel and scalable deep learning model for multi-event survival analysis, that learns the $K$ event distributions jointly, where each event is a convex combination of Weibull distributions. This approach relaxes the proportional hazards assumption about the underlying marginal survival distribution for each event, which is a significant limitation of Cox-based methods [6, 23, 33]. As in previous works [30, 51], we adopt a shared feature (covariate) layer in our architecture that captures the latent feature representation common to all events. Moreover, beyond simply optimizing the negative log-likelihood (NLL) over all events [51], we propose optimizing the sum of the NLL and a novel trajectory-based likelihood, which encourages the model to match the known temporal orderings between events based on prior knowledge. For example, consider the illness-death scenario [11] with two events “illness (A)” and “death (B)”: If both events are observed for the patient $i$, then clearly event A must occur before event B, and so we want to maximize the probability of surviving event B, at the time of event A – so if A happened at 1 year, we want the probability of surviving B at 1 year to be very high. MENSA is inspired by the multi-state framework, originally proposed by Andersen and Keiding [2], where an event can be considered a transition from one state to another – e.g., alive to dead, or healthy to sick. This framework enables us to model the event-free distribution (i.e., the transient state) explicitly as if it were another event and perform likelihood estimation in conjunction with the events of interest. This allows practitioners to predict whether a patient will remain event-free for the duration of the study, along with other events. In addition, the MENSA framework easily extends to the single-event and competing-risk scenarios, making it suitable for a wide range of applications.
@@
We provide experimental evidence to show that MENSA can accurately discriminate (i.e., rank) instances by risk and make time-to-event predictions that improve over several shallow and deep learning baselines. In particular, our model (1) relaxes the proportional hazard assumption, (2) improves over well-accepted baselines in several performance metrics (Harrell’s concordance index, Brier Score and MAE), and finally, (3) uses significantly fewer parameters and floating point operations per second (FLOPs) in high-dimensional datasets than many existing approaches. Ablation experiments demonstrate that applying a shared feature layer can improve predictive performances if conditional dependencies exist between events, and that the proposed trajectory likelihood improves the accuracy of the time-to-event estimates.
Our code is publicly available at https://github.com/thecml/mensa.
@@ 2 Related work
Single-event analysis. Time-to-event outcomes have been extensively studied within the field of survival analysis, primarily by the statistical, medical, and reliability engineering communities [5]. Traditional single-event models aim to learn the time until a single event occurs for an instance, for example, the time to death in amyotrophic lateral sclerosis (ALS) [28], the time to fall in older adults [32], or various other events in economic and social sciences. In practice, “survival prediction” does not necessarily refer to actual mortality. The challenge in survival prediction is that the event times can be partially observed (i.e., censored), providing only a lower bound on the true event time. Multiple censoring mechanisms may exist in practice, but in this work we will focus on the most common form: right-censoring, where we know the event occurs after a certain time, but not exactly when.
Figure 2: Illustration of right-censoring in survival datasets. Patients 1 and 2 experience the event. Patient 3 is lost to follow-up. Patient 4 withdraws. Patient 5 remains alive at the end of the study.
@@
In a medical study as shown in Figure 2, some patients may experience the event (1st and 2nd patient), but some may be lost to follow-up (3rd patient), withdraw from the study (4th patient), or remain alive at the end of the study (5th patient). For these patients, the exact event time is known only to exceed their last contact or the end of the study. A wide range of models have been proposed for this setting, including the nonparametric Kaplan-Meier (KM) [22] and Nelson-Aalen [37, 1] population estimators, nonparametric neural estimators for individuals, e.g., UniSurv [57], semiparametric proportional hazards models, e.g., CoxPH [6, 9, 23], and parametric survival models, e.g., the Weibull time-to-event model [38]. Chen [5] provide an excellent introduction to the field and describe the transition from classical time-to-event methods to modern deep learning approaches.
@@
Competing-risks analysis. In competing-risks survival analysis, there are multiple events of interest, e.g., for a patient, we consider the patient dying from lung cancer, or from pneumonia, as two different events, which are mutually exclusive. We thus want to learn a model that can predict, for each individual, which event will occur first and the time until this earliest event. Concerning censoring in such scenarios, competing risks is a form of dependent censoring [8, Ch. 1], as one risk (event) dependently censors the other, and these risks are assumed to have some prior relationship. However, estimating competing-risks data is more challenging than single-event due to the identifiability issue [52]; since events are mutually exclusive, the first event will censor the other events, preventing direct observation of their dependence structure. Given these challenges, multiple methods have been proposed for competing-risks scenarios, usually assuming that events are conditionally independent given the features.
@@
Lee et al. [30] proposed DeepHit, a discrete-time neural network of the form $f_{\theta}:\bm{x}\rightarrow[0,1]^{T_{\text{max}}\times K}$ with parameter $\theta$, number of events $K$, and some maximum observed time $T_{\text{max}}$. DeepHit contains a shared covariate layer and $K$ cause-specific sub-networks; its learning process learns $\theta$ by maximizing the log-likelihood of the joint distribution of $K$ competing events and a cause-specific ranking loss.
@@
Kim et al. [24] proposed MTLR-CR, an extension of the discrete Multi-Task Logistic Regression (MTLR) framework [55], a discrete-time survival model that predicts the survival distribution using a sequence of dependent logistic regressors. This is then extended to competing-risks scenarios by learning sequences of logistic regressors for the number of discrete time times, for every $K$ event.
@@
Nagpal et al. [36] proposed Deep Survival Machines (DSM), a parametric survival model based on a mixture of predefined primitive distributions, e.g., a mixture of Weibulls. DSM first learns a common feature representation of all competing risks, and then a $k$th-event distribution for each event using cause-specific maximum likelihood estimation (MLE), treating all events other than $k$ as censored. However, this approach does not explicitly model the density or survival function relationships between events on a one-to-one basis. As a result, it may not account for potential interdependencies between events.
@@
Multi-event analysis. In multi-event analysis, we again consider multiple critical events, but these are not necessarily mutually exclusive, as a patient may experience none, one or any subset of events at different times. For example, Mr. Smith may experience the onset of lung cancer at 1 year and may also experience the onset of pneumonia at 1.3 years, meaning that he would have both events (onset of lung cancer and of pneumonia). Multi-event analysis goes back as early as 1951, when Fix and Neyman [11] introduced the illness-death model, which has been used in multiple subsequent works [50]. Unlike competing-risks, multi-event analysis does not necessarily suffer from the identifiability issue [52], as the events are not necessarily mutually exclusive. However, predicting the outcome of several (possibly related) events instead of just one poses another challenge: When events are, in fact, competing, naive survival models that do not account for the dependency structure between events may produce biased estimates. Larson [29] adopted the popular CoxPH model [6] for multi-event survival analysis by modeling each event separately – an idea that has gained popularity since then [3, 49], but which fails to capture any dependencies between the events.
@@
Andersen and Keiding [2] proposed a multi-state model that learns a model for each event transition (e.g., given a person is diagnosed with lung cancer, it then predicts the probability of pneumonia), but this also fails to leverage shared covariate information among the events. Several works have addressed this restriction, e.g., using a shared frailty term across models [20], applying forms of regularization [31, 53] or by modeling the joint survival distribution [18]. However, these approaches require assumptions on the distribution of the frailty term, the proportional hazards assumption, or assumptions on the form of the joint survival distribution, respectively.
@@
Tjandra et al. [51] proposed a hierarchical multi-event model that learns the probability of event occurrence at different hierarchical time scales, using coarse predictions to iteratively guide predictions at finer levels – for example, from monthly to daily predictions. Given a complex task over a long time span, the model decomposes this task into a sequence of simpler tasks that serve as intermediate guides. Thus, the hierarchical framework progressively predicts survival across multiple time granularities, starting with broader intervals (e.g., months) and refining down to finer ones (e.g., weeks). Formally, their model is a multi-task network that takes $x$ as input and outputs $\hat{P}$, a set of probability distributions for each $k$ event. Event-specific subnetworks, $\phi_{k}$, estimate the probability distribution $\hat{p}_{k}$ hierarchically. At granularity $m$, the probability distribution $\hat{p}_{km}$ is constructed from $\psi_{km}(\Phi(x))$ and the previous granularity level $\hat{p}_{k,m-1}$, ensuring that predictions at finer time scales are guided by coarser ones. There are two primary limitations to this approach: (1) the hierarchical component scales poorly with larger time durations and has a significant parametric overhead. (2) the method is optimized solely for discriminative loss –- i.e., ranking using Harrell’s concordance index (CI) [17]. However, strong ranking performance does not necessarily translate into accurate time-to-event predictions [40], and vice versa. In our experiments, we highlight these two limitations to the hierarchical model, i.e., its computational overhead and poor predictive accuracy.